OpenStudy (anonymous):

approximate the area under the curve x^2-2 from -2 to1 using RRAM

OpenStudy (amistre64):

define what RRAM is

OpenStudy (anonymous):

right Reiman approximation method is what it is and the question needs to ask you how many segments it is broken into.

OpenStudy (amistre64):

"approximate" lets you use your own division of the interval, it all depends on how approx you want to make it

OpenStudy (amistre64):

\[\sum_{i=0}^{n-1}f(a+i\frac{b-a}{n})\frac{b-a}{n}\] \[\sum_{i=0}^{n-1} ((-2+i\frac{1+2}{n})^2-2)\frac{1+2}{n}\] \[\sum_{i=0}^{n-1} ((-2+\frac{3i}{n})^2-2)\frac{3}{n}\] pick your favorite n :)